OpenStudy (anonymous):

-16x^2+9y^2-160x+12y-400=0 need to graph

OpenStudy (anonymous):

Find the tables of values with values of x and corresponding values of y for example when x=1 , y=something,.......and plot in graph......

OpenStudy (anonymous):

complete da square

OpenStudy (anonymous):

-16(x^2+10x) + 9( y^2 + 4/3 y ) = 400

OpenStudy (anonymous):

-16(x^2 + 10x + 25 - 25) + 9(y^2 +4/3y + 16/36 - 16/36) = 400

OpenStudy (anonymous):

-16(x+5)^2 + 25*16 + 9(y+2/3)^2 - 9*16/36 = 400

OpenStudy (anonymous):

-16(x+5)^2 + 9(y+2/3)^2 = 400 - 25*16 + 9*16/36 = 4

OpenStudy (anonymous):

(x+5)^2/(1/2)^2 + (y+2/3)^2 / (2/3)^2 = 1

OpenStudy (anonymous):

an ellipse centered at (-5, -2/3) with a width of 1/2 and a height of 2/3

OpenStudy (anonymous):

missed a sign in that last step, it's -(x+5)^2/(1/2)^2 + (y+2/3)^2 / (2/3)^2 = 1

OpenStudy (anonymous):

a hyperbola